# Towards Open-World Gesture Recognition

### v1

@@ 7.1.2 How to update the recognition model?
We observe that naive finetuning experiences serious catastrophic forgetting and is not suitable for OWGR. For the new context case, the replay method outperforms other continual learning methods by a large margin in accuracy. In contrast, PackNet has zero forgetting by its design and still maintains relatively high accuracy. Overall, for the new context case, the replay method is the best choice to optimize accuracy – memory/privacy constraints permitting – whereas PackNet is a better choice to balance accuracy and forgetting, provided the model capacity is sufficient for the total number of tasks.
For new gesture, the replay method does not show a significant advantage in accuracy over other continual learning methods. In contrast, both PackNet and LwF yield a high accuracy with low forgetting. Therefore, LwF is a better choice overall for the new gesture case, as unlike PackNet, it does not impose constraints on total model size.
For new user, the replay method no longer ranked as the top performer in accuracy. Both PackNet and LwF exhibit more stable performance in this case as the difference between the first quartile (Q1) and the third quartile (Q3) is smaller, showing a smaller variance in the performance metrics. Thus PackNet and LwF is a better choice for the new user case.
@@
7.2 Envelope Analysis: A Practical Alternative to Large-Scale User Studies
This research aims to present a novel approach for evaluating continuous learning methods in open-world gesture recognition, eliminating the need for extensive user studies. We propose envelope analysis combined with design engineering as a practical alternative to traditional user studies. This method allows us to simulate various scenarios and environmental conditions, which typically require large-scale user studies. Our confidence in this approach is backed by an extensive and diverse dataset, reducing the need for separate user studies.
@@ 7.3 Limitation and Future Work
One limitation of our study is the lower accuracy in gesture recognition. It’s worth noting that the accuracy reported in this paper represents the final model’s accuracy on the last task, following the entire task sequence (see Section 5.3). The accuracy diminished from the initial task to the final task due to catastrophic forgetting. This further encourages us to refine our continual learning algorithms as future work. Another contributing factor is that our training algorithm wasn’t explicitly tailored for optimal accuracy. Techniques such as data augmentation, feature engineering, and model pre-training, which are commonly used to enhance model performance [63, 49, 47], were not employed in our study. This decision was intentional, as our primary aim was to isolate the effects of the envelope analysis. By excluding these optimization techniques, we ensured that any observed effects on the final results were solely due to changes in the parameters of the envelope analysis.
@@ 8 Conclusion
This paper presents a design engineering approach, using a large-scale dataset and envelope analysis, to explore continual learning methods for the Open-World Gesture Recognition (OWGR) process. This approach is applicable even when training and testing data do not share similar characteristics, encompassing new contexts, new users, and new gestures. We discuss the significant catastrophic forgetting observed in the baseline finetuning approach and examine the pros and cons of various continual learning methods. The paper’s main contribution lies in presenting these guidelines derived from extensive data collection and computational experiments. Intended for developers and designers of wrist-worn gesture-sensing systems, our guidelines provide an alternative to traditional large-scale user studies, enhancing accessibility and promoting a more refined OWGR process. Our design engineering methodology, backed by envelope analysis, offers a new route for system evaluations, potentially inspiring others to adopt similar methods for real-world assessment challenges.

### v2

@@ 5.2.2 How do we update the recognition model?
We observe that naive fine-tuning experiences serious catastrophic forgetting and is not suitable for OWGR. For the new context case, the replay method outperforms other continual learning methods by a large margin in accuracy. In contrast, PackNet has zero forgetting by its design and still maintains relatively high accuracy. Overall, for the new context case, the replay method is the best choice to optimize accuracy—memory/privacy constraints permitting—whereas PackNet is a better choice to balance accuracy and forgetting, provided the model capacity is sufficient for the total number of tasks. For new user, the replay method no longer ranked as the top performer in accuracy. Both PackNet and LwF exhibit more stable performance in this case as the difference between the first quartile (Q1) and the third quartile (Q3) is smaller, showing a smaller variance in the performance metrics. Thus PackNet and LwF is a better choice for the new user case.
@@
5.3 Envelope Analysis: A Practical Alternative to Large-Scale User Studies in OWGR
The primary objective of this research is to introduce an novel method for assessing continual learning methods for OWGR, bypassing the need for large-scale user studies. Researchers often lack resources for such studies, which are more demanding than evaluating conventional gesture recognition models in controlled settings. The proposed approach uses envelope analysis with parameterization of key functions, offering a robust and practical alternative to user studies. By varying parameters, different scenarios can be systematically explored. The extensive and diverse dataset used in this research provides a strong foundation for evaluating the OWGR process, reducing the need for a separate large-scale user study. Although some conclusions drawn from the envelope analysis might appear intuitive, there has been a notable absence of such thorough experiments in the realm of what we would classify as OWGR. Further, assumptions from other, related, areas might not directly translate to gesture recognition problems relevant for mixed reality applications.
5.4 OWGR in Mixed Reality
OWGR enables gesture recognition models to adaptively learn new gesture data patterns from diverse contexts and users on-the-fly, eliminating the need for extensive pre-deployment data collection. It facilitates personalization by learning and adapting to individual users’ contexts, whether performed in static or dynamic environments. For shared devices among multiple users, such as families, it eliminates recalibration needs by remembering previous users’ patterns while continuously learning from new user. This ability to seamlessly learn and retain knowledge across contexts and users ensures a robust, calibration-free experience.
@@ 5.5 Limitation and Future Work
One limitation in our study is the lower accuracy in gesture recognition. It’s worth noting that the accuracy reported in this paper represents the final model’s accuracy on the last task, following the entire task sequence (see Section 3.3). The accuracy diminished from the initial task to the final task due to catastrophic forgetting. This further encourages us to refine our continual learning algorithms as future work. Another contributing factor is that our training algorithm wasn’t explicitly tailored for optimal accuracy. Techniques such as data augmentation, feature engineering, and model pre-training, which are commonly used to enhance model performance [66, 51, 49], were not employed in our study. This decision was intentional, as our primary aim was to isolate the effects of the envelope analysis. By excluding these optimization techniques, we ensured that any observed effects on the final results were solely due to changes in the parameters of the envelope analysis.
@@ 6 Conclusion
This paper presents a design engineering approach, using a large-scale dataset in combination with envelope analysis, to explore continual learning methods for the open-world gesture recognition process. We discuss the significant catastrophic forgetting observed in the baseline fine-tuning approach and examine the positive and negative qualities of various continual learning methods. To assist other researchers and developers we present guidelines derived from the extensive data collection and the computational experiments. Intended for developers and designers of wrist-worn gesture-sensing systems, our guidelines provide an alternative to traditional large-scale user studies, lowering the barriers and promoting a more refined OWGR process. This paper is also a demonstration of how a design engineering methodology, in particular envelope analysis, offers a new route for system evaluations in this area, potentially inspiring other researchers in mixed reality to adopt similar methods for real-world assessment challenges.